Can you help me solve this problem. A container holds 3 liters of a 25 % alcohol solution. One half liter of the solution is withdrawn and replaced with one half liter of pure alcohol. What is the percentage of alcohol in new mixture ?
Answer by mananth(16946)

You are adding 1.5 liters of 100% alcohol to 25% solution to get 3 liters of mixture.
Let the percentage of mixture be x
1.5*100% + 1.5*25% = 3*x%
multiply by 100
1.5+37.5=3x
39=3x
x=13 %